OpenStudy (anonymous):

Solve the equation 3 over/ 7x + 5 = 8

sammixboo (sammixboo):

\[\frac{ 3 }{ 7x }+5=8\] Is that how your equation looks like?

OpenStudy (anonymous):

\[\frac{ 3}{ 7} x +5 = 8\] @sammixboo

OpenStudy (anonymous):

\[A. 1\frac{ 2 }{ 7 } b. 7\frac{ 2 }{ 3 } c. -7 d. 7\]

OpenStudy (phi):

I would add -5 to both sides as a first step. can you do that ?

sammixboo (sammixboo):

1st you do what @phi says... Subtract 5 on both sides of the equations 2nd you divide by 3 on both sides of the equation. 3rd You multiply 7 on both sides... Whatever is on the RIGHT side of the equation is your answer...

OpenStudy (anonymous):

is the answer 7 @sammixboo

sammixboo (sammixboo):

Yes! Yay! I am glad I could help you!!

sammixboo (sammixboo):

Do you got any other questions?

OpenStudy (anonymous):

yes @sammixboo last one \[\frac{ 2}{ 3}x -6 = 9\]

sammixboo (sammixboo):

Ok.... 1st you add 6 to both sides 2nd divide 2 on both sides 3rd you multiply 3 on both sides...

OpenStudy (anonymous):

is it 10 @sammixboo

sammixboo (sammixboo):

Do you have any choices???

OpenStudy (anonymous):

yes 10, \[22\frac{ 1 }{ 2 } , -22 \frac{ 1}{ 2 } and 25\]

OpenStudy (anonymous):

@sammixboo

sammixboo (sammixboo):

No.. It is \[22 \frac{ 1 }{ 2 }\] Because: 1: \[\frac{ 2x }{ 3 } - 6 = 9 \] 2: \[\frac{ 2x }{ 3 } - 6 + 6 = 9 + 6\] 3: \[\frac{ 2x }{ 3 } = 15\] 4: \[\frac{ 2x \div 2 }{ 3 } = 15 \div 2\] 5: \[\frac{ x }{ 3 } = 7 \frac{ 1 }{ 2 }\] 6: \[3 \times \frac{ x }{ 3 } = 7 \frac{ 1 }{ 2 } \times 3\] 7: \[x = 22 \frac{ 1 }{ 2 }\]

sammixboo (sammixboo):

Wow that took a while *wipes sweat off brow* Whew

sammixboo (sammixboo):

4 minutes...

OpenStudy (anonymous):

ok great thanks for your help :) @sammixboo

sammixboo (sammixboo):

You are welcome! Just tag me in any question and I will help!

sammixboo (sammixboo):

Thanks @phi for helping me remember on how to do this!